OpenStudy (anonymous):

At the corner diner, 3 pretzels and 1 cup of soda cost $2.75. Two pretzels and 1 cup of soda cost $2.00. Find the cost of a pretzel and the cost of a cup of soda. Help me!!!!! the topic is the solving the Systems of Equations.

OpenStudy (anonymous):

these are really neat! Let us try and call pretzels x and soda y. can you write this out for me? 3 pretzels and 1 cup of soda costs 2.75. use x and y and turn that into an equation please!

OpenStudy (anonymous):

I think it will easier to make the pretzel a (p) and the soda(s). so then there are two equations. First one will be 3p+s=2.75 and the second will be 2p+s=2.00.

OpenStudy (anonymous):

that works too! now, pick one of those equations (i like the second 1 because 2 is a nice smooth number but either is fine) and tell me what s equals in terms of p. (just get s alone on 1 side)

OpenStudy (anonymous):

i cant do it anymore. Its already so hard!!!!!

OpenStudy (anonymous):

ok i got ya. you are already 1/3 the way done though

OpenStudy (anonymous):

equation 2. 2p + s = 2.00 subtract 2p from each side! s = 2.00 - 2p how does this look to ya?

OpenStudy (anonymous):

we are only worried about equation 2 right now! baby steps for this one.

OpenStudy (anonymous):

oh, so then i substitute and plug in the numbers right?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

i personally like to put ( ) around the s's first.

OpenStudy (anonymous):

then put inside of the the parenthesis, whatever s equals. this will keep you from mixing things up!

OpenStudy (anonymous):

3p+s=2.75 3p+(2.00-2p)=2.75 1p+2.00=2.75 subtract 2.00 from both sides 1p=.75

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

so now I know that one cup of soda costs 75 cents

OpenStudy (anonymous):

well, p is pretzel heheh

OpenStudy (anonymous):

so, now you plug in for the p and you solve for s! then you are done!

OpenStudy (anonymous):

then I find about the pretzels. 3p+s=2.75 3p+.75=2.75 subtract .75 from both sides 3p=2.00 divide it by 3 from both side then get .66

OpenStudy (anonymous):

wait, i dont get the dividing part at the end

OpenStudy (anonymous):

oops! you have 1p = .75 right? that means... 1p(retzel)=.75!

OpenStudy (anonymous):

does that mean each pretzel cost 66 cents?

OpenStudy (anonymous):

no no. p = .75 means 1 pretzel = .75! so... every spot you see a p put it in ( ) and put .75 in there!

OpenStudy (anonymous):

3p+s=2.75 3(p) + s = 2.75 3(.75) + s = 2.75 3*.75 = 2.25 sooooooo 2.25 + s = 2.75

OpenStudy (anonymous):

ohhhhhhh.... so then you minus 2.25 from both sides and then get s=50 cents right?

OpenStudy (anonymous):

so... one pretzel cost 75 cents and one cup of soda costs 50 cents.

OpenStudy (anonymous):

it seems a bit confusing because first you think you are solving for s right? because you got s = 2.00-2p. what is ACTUALLY going on is... you are getting rid of the s so that you only have one variable.

OpenStudy (anonymous):

yes! remember! if you see 1p = .75 then you literally solved the cost of 1 pretzel! 1 pretzel = 75 cents!

OpenStudy (anonymous):

kk. Thank you so much for the help!! Really appreciated it!! good night. :)